OpenStudy (anonymous):

Find the slope of the line that contains the following points. A(5, 6), B(10, 8)

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

hey bri

OpenStudy (amistre64):

move the points so that one of them is at the origin .... then the slope is easier to see

OpenStudy (amistre64):

or hack out the slope formula ....

terenzreignz (terenzreignz):

Slope is all about differences :) Literally, the difference of the y's divided by the difference of the x'x ^_^

OpenStudy (austinl):

\[m = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (amistre64):

A(5, 6), B(10, 8) 5 -6 -5 -6 -------------- 0 0 5 2

terenzreignz (terenzreignz):

^neat :)

OpenStudy (austinl):

\[m = \frac{8-6}{10-5}=\frac{2}{5}\]

OpenStudy (anonymous):

Thanks @austinL

OpenStudy (austinl):

Your method is pretty neat @amistre64

terenzreignz (terenzreignz):

Forgot a minus-sign, from the looks of it :3

OpenStudy (amistre64):

thnx .. i tend to misplace the parts into the formula so i just step it thru

OpenStudy (amistre64):

i dint forget ... i went to move that line and it deleted lol